Solve:
(a+b)^3
Thanks

Found 2 solutions by checkley71, funmath:
Answer by checkley71(8403)

(A+B)^3
(A+B)(A+B)=(A^2+2AB+B^2)(A+B)=A^3+2A^2B+AB^2+A^2B+2AB^2+B^3=
A^3+3A^2B+3AB^2+B^3

Answer by funmath(2933)

Solve:
(a+b)^3
Most people end up having to memorize this, this is how you derive it.
(a+b)(a+b)(a+b) Multiply the second two.
(a+b)(a^2+ab+ab+b^2)
(a+b)(a^2+2ab+b^2) distribute the a then the b from the first parentheses.
a(a^2+2ab+b^2)+b(a^2+2ab+b^2)
a^3+2a^2b+ab^2+a^2b+2ab^2+b^3
a^3+2a^2b+a^2b+ab^2+2ab^2+b^3
a^3+3a^2b+3ab^2+b^3
If you're in algebra II it's worth memorizing.
Happy Calculating!!!